For this graphical display of Quiz Scores, which estimates of the mean and median are most plausible? 3 2 6 10 14 18 22 uiz Seores median = 13.0 and mean = 12.0 median 14.0 and mean 15.0 median16.0 and mean 14.3 median = 16.5 and mean = 16.2
Answer #1

Given is a graphical display of Quiz Scores, now we need to find which estimates of the mean and median are most plausible.

Now clearly the the graphical display is a histogram with classes of length 2 in the x axis and the corresponding frequencies in the y axis. Now let us re-make the frequency distribution which corresponds to the histogram.

No. Class Boundary Frequency
Lower Upper
1 4 6 1
2 6 8 1
3 8 10 0
4 10 12 1
5 12 14 2
6 14 16 3
7 16 18 5
8 18 20 4
Total 17

Now let us calculate the median of this frequency distribution. For that we prepare a cumulative frequency column of less than type.

No. Class Boundary Frequency Cumulative Frequency
Lower Upper
1 4 6 1 1
2 6 8 1 2
3 8 10 0 2
4 10 12 1 3
5 12 14 2 5
6 14 16 3 8
7 16 18 5 13
8 18 20 4 17
Total 17

Now median is given by,

Median = x + n/2-ni

01 = Lower Class Boundary of the class containing the median

n = Total frequency

ni = Cumulative frequency of the previous class containing the median

fo = Frequency of the class containing the median

c=\emph{Class Width}

Now here the median class is, n/2=17/2=8.5 i.e the 7th class is the median class.

Hence,

x_{l}=16,n=17,n_{l}=8,f_{0}=5,c=2

\Rightarrow Median=16+\frac{17/2-8}{5}*2

\Rightarrow Median=16+0.2

\Rightarrow Median=16.2

Now mean is given by,

Mean=\frac{1}{n}\sum_{i=1}^{k}x_{i}f_{i}

where,

n=\sum_{i=1}^{k}f_{i}

x_{i}=\emph{Class Mark of the ith class}

f_{i}=\emph{Frequency of the ith class}

k=\emph{Number of classes}

So we make a class mark column our frequency distribution. Now class mark is the average of the upper and class boundary.

No. Class Boundary Class Mark(xi) Frequency(fi) xi*fi
Lower Upper
1 4 6 5 1 5
2 6 8 7 1 7
3 8 10 9 0 0
4 10 12 11 1 11
5 12 14 13 2 26
6 14 16 15 3 45
7 16 18 17 5 85
8 18 20 19 4 76
Total 17 255

=*==0.

\Rightarrow Mean=15

Hence Median=16.2 and Mean=15.

Hence the most plausible option is third option Median=16 and Mean=14.3. Since they are close to the exact mean and median.